If the polynomials x^3+ax^2+3x+5x and x^3 + 2x^2 + x + 2a leave the same remainder when divided by (x−1), then the value of a is:


1
Expert's answer
2021-10-01T12:55:58-0400

"x-1=0=>x=1"



"(1)^3+a(1)^2+3(1)+5=(1)^3+2(1)^2+1+2a"

"a+9=2a+4"

"a=5"